OpenStudy (anonymous):

A floor has two square-shaped designs. The area of the second square-shaped design is twenty-five times greater than the area of the first square-shaped design. Which statement gives the correct relationship between the lengths of the sides of the two squares? Answer The length of the side of the second square is 25 times greater than the length of the side of the first square. The length of the side of the second square is 5 times greater than the length of the side of the first square. The length of the side of the second square is 50 times greater than the length of the side

OpenStudy (experimentx):

lt the length of one square be x, then it's area is x*x = x^2 according to the question, the length of other square is 25* x^2 = (5x)^2 the it's length is 5x I hope you got your answer

OpenStudy (anonymous):

ok yes thank -you the next question i have on here is similar but i think i have the answer The sun roof in Troy’s house is shaped like a rectangle. He increases the length of the sides of the sun roof to five times the existing ones. How will this change affect the perimeter of the sun roof? Answer It will be 10 times the original perimeter. It will be 15 times the original perimeter. It will be 20 times the original perimeter. It will be 5 times the original perimeter. i think it's five right

OpenStudy (experimentx):

perimeter = x + y + x + y new perimeter = x + y+5 + x + y + 5

OpenStudy (anonymous):

ok thanks